Question 3: Ibus Ltd is considering the installation of a new computer. Because of uncertainty as to its future computing requirements and the prospect of advancements in computing technology, it is evaluating the acquisition of the computer by either purchasing it, or leasing it under a contract that includes a cancellation option. Information relevant to the company's evaluation is as follows: Purchase The purchase price of the computer is $300 000 and it can be depreciated at a rate of 15 per cent per annum, straight-line. Ibus Ltd plans to operate the computer for a maximum of 5 years. The computer's disposal value at the end of 5 years is estimated to be $50 000. Lease The annual lease payments on the operating lease would be $90 000, payable at the beginning of each year. The lease can be cancelled by Ibus Ltd at any time without incurring any penalty payment. The company income tax rate is 30 cents in the dollar, the required return on the investment is 20 per cent per annum and the after-tax cost of an equivalent loan is 10 per cent per annum. Should the company purchase or lease the asset?
Answer #1
We shall use the after tax cost of Equivalent loan , not return on
investment as lease/buy decision generally does not involve the
operating risk.
After Tax cost of Loan =10% pa
Option 1. Purchases
Purchase cost $               300,000
Depreciation per annum @15%= $            45,000.00
Total depreciation in 5 yrs= $         225,000.00
Written down value after 5 yrs= $            75,000.00
Disposal value after 5 yrs $            50,000.00
Capital Loss on Disposal $            25,000.00
Tax savings on capital loss@30%= $              7,500.00
Salvage value+Tax savings= $            57,500.00
Finding NPV of Purchase a b c d e f
Year Initial Investment Annual Depreciation Depreciation Tax Shield @30 Salvage value with tax savings Net Cash flow =a+c+d PV factor @10%=1/1.1^n PV of Net cash flow=e*f
0 $             (300,000) $       (300,000) 1 $    (300,000)
1 $              45,000 $                 13,500.0 $        13,500.0 0.9091 $        12,273
2 $              45,000 $                 13,500.0 $        13,500.0 0.8264 $        11,156
3 $              45,000 $                 13,500.0 $        13,500.0 0.7513 $        10,143
4 $              45,000 $                 13,500.0 $        13,500.0 0.6830 $          9,221
5 $              45,000 $                 13,500.0 $         57,500.00 $        71,000.0 0.6209 $        44,084
NPV = $    (213,124)
Option : Operating Lease a. b c d e
Year Annual Lease Tax shield @30% on lease rental Net After Tax cost PV factor @10%=1/1.1^n PV of Net cash flow=c*d=
1 $               (90,000) $          27,000.0 $               (63,000.0) 0.9091 $    (57,273.30)
2 $               (90,000) $          27,000.0 $               (63,000.0) 0.8264 $    (52,063.20)
3 $               (90,000) $          27,000.0 $               (63,000.0) 0.7513 $    (47,331.90)
4 $               (90,000) $          27,000.0 $               (63,000.0) 0.6830 $    (43,029.00)
5 $               (90,000) $          27,000.0 $               (63,000.0) 0.6209 $    (39,116.70)
NPV $ (238,814.10)
So NPV of purchase is higher than NPV of lease.
Therfore, the company should purchase the asset.